OpenStudy (anonymous):

A person has 20 coins, all nickels and dimes, worth $1.65. How many nickels are there? 7 8 13

Parth (parthkohli):

If the man has x nickels, then he has 20 - x dimes.

OpenStudy (anonymous):

ok

Parth (parthkohli):

Nickels are 5 cents. Dimes are 10 cents. He has 165 cents. \( \color{Black}{\Rightarrow x(5) + (20 - x)(10) = 165 }\)

OpenStudy (anonymous):

ok so what do i solve first in this one

Parth (parthkohli):

Well you must distribute the values, first of all.

Parth (parthkohli):

Do you know how to distribute or afk?

OpenStudy (anonymous):

ok so ya

OpenStudy (anonymous):

kinda

Parth (parthkohli):

\( \color{Black}{\Rightarrow x(5) = ....? }\)

OpenStudy (anonymous):

5x

Parth (parthkohli):

\( \color{Black}{\Rightarrow 10(20 - x) = ....? }\)

OpenStudy (anonymous):

30-x

Parth (parthkohli):

No, I'm not sure what you did here, but: \( \color{Black}{\Rightarrow a(b \pm c) = (a \times b) \pm (a \times c) }\)

OpenStudy (anonymous):

wouldnt it be 200-x

Parth (parthkohli):

This means: \( \color{Black}{\Rightarrow a(b + c) = (a \times b) + (a \times c) }\) OR \( \color{Black}{\Rightarrow a(b - c) = (a \times b) - (a \times c) }\)

Parth (parthkohli):

200 is correct....but what about \(10 \times -x\)?

OpenStudy (anonymous):

-10x

Parth (parthkohli):

Correct. So back to the equation. \( \color{Black}{\Rightarrow 5x + (200 - 10x) = 165 }\)

OpenStudy (anonymous):

ok now get rid of the brackets

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

-200 right

Parth (parthkohli):

Well if there is a positive sign outside the brackets, then you just remove them like *SWOOSH*

Parth (parthkohli):

If there is a negative sign, then you change the signs of the stuff in the bracket.

OpenStudy (anonymous):

ok so then i just solve like *SWOOSH*

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5x + 200 - 10x = 165 }\) I just *SWOOSH*ed them out ;)

OpenStudy (anonymous):

combine like terms

Parth (parthkohli):

Yes, what do you get?

OpenStudy (anonymous):

-5x

Parth (parthkohli):

\( \color{Black}{\Rightarrow -5x + 200 = 165 }\) What do we do now?

OpenStudy (anonymous):

-5x=-35

Parth (parthkohli):

Yes, what do we do now?????????

OpenStudy (anonymous):

x=-7

Parth (parthkohli):

No....! you are very very close.

Parth (parthkohli):

\( \color{Black}{\Rightarrow -a = -b \Longleftrightarrow a = b }\)

OpenStudy (anonymous):

double b\negative right so 7

OpenStudy (lgbasallote):

side note: if you think about it negative number of money doesnt make sense ;D

Parth (parthkohli):

yes. 7.

OpenStudy (anonymous):

ok thanks parth

Parth (parthkohli):

Pleasure :D